A thin layer of oil (n = 1.50) floats on water (n = 1.33). Destructive interference is observed for light of wavelength 470 nm. Find the minimum non-zero thickness of the oil film assuming normal incidence.
Answer #1

n = 1.33, λ = 470 nm, find tmin

Note 1.50 > 1.33 > 1, the path difference = 2nt = kλ

t = kλ/(2n)

tmin = λ/(2n) = 470/(2*1.33) = 176.70 nm----------Answer